# Misiurewicz points and subhyperbolicity in unicritical algebraic correspondences

### v1

@@ 7. Sub-hyperbolicity
There exists a simply connected set $U_{1}$ containing $w_{0}$ and $\tilde{z}_{0}$ such that $U_{1}$ is contained in the complement of the set of ramified points $R.$ Therefore, we are allowed to compute the length $\ell_{\rho}(\gamma)$ of curves in $U_{1}$ with respect to the orbifold metric $\rho$, as well as the diameter $|U_{1}|_{\rho}$ of $U_{1}$ with respect to $\rho.$ Choose some closed ball $\tilde{K}$ of large radius whose complement is a forward invariant set contained in the basin of infinity $\mathbb{C}{\setminus}K_{a}.$ Hence $\mathbf{f}_{a}^{-1}(\tilde{K})\subset\tilde{K}$ and we may assume that $K_{a}$ is contained in the interior of $\tilde{K}$. Since the set of ramified points is locally finite, $\tilde{K}$ contains only finitely many points of $R;$ it is not hard to check that we can always take $U_{1}\subset\tilde{K}$ satisfying the following property: there exists a positive constant $C_{0}$ such that any two points of $U_{1}$ are joined by a curve $\gamma$ within $U_{1}$ with $\ell_{\rho}({\gamma})\leq C_{0}$ (indeed, it suffices to take $U_{1}$ as an $\epsilon$-neighborhood of some curve joining $w_{0}$ and $\tilde{z}_{0}$ avoiding the finite set $\tilde{K}\cap R$).
Since $U_{1}$ is simply connected and the critical value $a$ is not in $U_{1}$, by analytic continuation there exists a unique branch of $\mathbf{f}_{a}^{-1}$ defined on $U_{1}$ which sends $\tilde{z}_{0}$ to $\tilde{z}_{1}$. The set $\tilde{K}{\setminus}R$ is backward invariant under $\mathbf{f}_{a}$ and contains $U_{1}$; thus $g_{1}(U_{1})$ is also contained in $\tilde{K}{\setminus}R.$ Inductively, we construct an infinite sequence of branches $g_{n}$ of $\mathbf{f}_{a}^{-n}$ such that $g_{n}$ is the unique branch of $\mathbf{f}_{a}^{-1}\circ g_{n-1}$ defined on $U_{1}$ which sends $\tilde{z}_{0}$ to $\tilde{z}_{n}.$
@@
(7.5) $\Omega_{1}{\setminus}\tilde{L}\xrightarrow{h^{-1}}W_{1}{\setminus}\tilde{L}_{a}\xrightarrow{F}V_{1}\xrightarrow{(f_{n}|_{U_{2}})^{-1}}U_{2}\xrightarrow{g_{k}}O_{2}\xrightarrow{f^{-n_{1}}}\Omega_{2}\Subset\Omega_{1}{\setminus}\tilde{L}$
@@ Acknowledgments
The author is grateful to Daniel Smania for his hospitality at ICMC/USP, and to Luna Lomonaco for helpful comments. Research partially supported by CNPq/MCTI/FNDCT, 406750/2021-1.

### v2

@@ 7. Subhyperbolicity
Let $U$ be an arbitrary conformal disk intersecting $K_{a}.$ By Lemma 7.2, a holomorphic branch $H$ of some iterate of the correspondence $\mathbf{f}_{a}$ sends $U$ to a connected neighborhood $V$ of zero, with $H^{\prime}(z)\neq 0$ for all $z\in U.$ Thus $H(w_{0})=0$ for some $w_{0}\in U$. Notice that $w_{0}\not\in R$, since $a$ is Misiurewicz. Moreover, $W=\mathbf{f}_{a}(V)$ is a connected neighborhood of $a$.
We may choose a simply connected region $U_{1}$ containing $w_{0}$ and $\tilde{z}_{0}$ such that $\overline{U_{1}}$ is contained in the complement of the set of ramified points $R.$
There exists a conformal map $h$ defined on a neighborhood of $a$ which is a branch of some iterate $\mathbf{f}_{a}^{k_{0}}$ and sends $a$ to $\zeta_{0}.$ By choosing $s\in(0,r)$ sufficiently small, we may assume that $h$ maps a neighborhood $W_{1}$ of $a$ contained in $W$ biholomorphically onto $\Omega_{1}=\varphi(\mathbb{D}_{s})\subset\Omega.$ Let $V_{1}$ denote $\mathbf{f}_{a}^{-1}(W_{1})$, which is a connected neighborhood of zero contained in $V.$ If we reduce the size of $s$, then both sets $V_{1}$ and $W_{1}$ will shrink accordingly; since $H$ is conformal and sends $w_{0}$ to zero, by reducing $s$ if necessary we may assume that $H$ sends a small connected neighborhood $U_{2}\subset U_{1}$ of $w_{0}$ biholomorphically onto $V_{1}.$ It is important to notice that $U_{2}$ also depends on $s$, and if $s\to 0$, then $U_{2}$ will shrink to $\{w_{0}\}.$
Since $U_{1}\subset\mathbb{C}{\setminus}R$, we are allowed to compute the length $\ell_{\rho}(\gamma)$ of curves in $U_{1}$ with respect to the orbifold metric $\rho$, as well as the diameter $|U_{1}|_{\rho}$ of $U_{1}$ with respect to $\rho.$ Choose some closed ball $\tilde{K}$ of large radius whose complement is a forward invariant set contained in the basin of infinity $\mathbb{C}{\setminus}K_{a}.$ Hence $\mathbf{f}_{a}^{-1}(\tilde{K})\subset\tilde{K}$ and we may assume that $K_{a}$ is contained in the interior of $\tilde{K}$. Since the set of ramified points is locally finite, $\tilde{K}$ contains only finitely many points of $R;$ it is not hard to check that we can always take $U_{1}\subset\tilde{K}$ satisfying the following property: there exists a positive constant $C_{0}$ such that any two points of $U_{1}$ are joined by a curve $\gamma$ within $U_{1}$ with $\ell_{\rho}({\gamma})\leq C_{0}$ (indeed, it suffices to take $U_{1}$ as an $\epsilon$-neighborhood of some curve joining $w_{0}$ and $\tilde{z}_{0}$ avoiding the finite set $\tilde{K}\cap R$).
Since $U_{1}$ is simply connected and the critical value $a$ is not in $U_{1}$, by analytic continuation there exists a unique branch $g_{1}$ of $\mathbf{f}_{a}^{-1}$ defined on $U_{1}$ which sends $\tilde{z}_{0}$ to $\tilde{z}_{1}$. The set $\tilde{K}{\setminus}R$ is backward invariant under $\mathbf{f}_{a}$ and contains $U_{1}$; thus $g_{1}(U_{1})$ is also contained in $\tilde{K}{\setminus}R.$ Inductively, we construct an infinite sequence of branches $g_{n}$ of $\mathbf{f}_{a}^{-n}$ such that $g_{n}$ is the unique branch of $\mathbf{f}_{a}^{-1}\circ g_{n-1}$ defined on $U_{1}$ which sends $\tilde{z}_{0}$ to $\tilde{z}_{n}.$
@@
(7.5) $\Omega_{1}{\setminus}\tilde{L}\xrightarrow{h^{-1}}W_{1}{\setminus}\tilde{L}_{a}\xrightarrow{F}V_{1}\xrightarrow{(H|_{U_{2}})^{-1}}U_{2}\xrightarrow{g_{k}}O_{2}\xrightarrow{f^{-n_{1}}}\Omega_{2}\Subset\Omega_{1}{\setminus}\tilde{L}$
@@ Acknowledgments
The author sincerely thanks Daniel Smania for his kind hospitality at ICMC/USP and Luna Lomonaco for valuable comments and insightful discussions. This work was partially supported by CNPq/MCTI/FNDCT under grant 406750/2021-1. The author is grateful to the anonymous referee for a careful reading of the manuscript and helpful suggestions that improved the presentation.